Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Coin Probab

Expert replies
by zagcollins » Sun Jul 20, 2008 5:26 am
When tossed, a certain coin has equal probability of landing on either side. If the coin is tossed 3 times, what is the probability that it will land on the same side each time?

A. 1/8
B. 1/4
C. 1/3
D. 3/8
E. 1/2

my answer is B...wrote down the probable outcomes..i.e.

HHH
TTT
HTT
HHT
THH
TTH
THT
HTH

Now we need heads on all three tries or tails on all three tries...so using the formula for discrete probab: desired outcome/total possible outcome

2/8= 1/4

I think my solution and approach, both, are correct...Pls advise if there's a simpler way to tackle this problem...
Join the discussion
Source: — Problem Solving |

by Saule » Sun Jul 20, 2008 7:42 am
I think it's actually 1/8 since 1/2*1/2*1/2=1/8.
Join the discussion

by Saule » Sun Jul 20, 2008 8:08 am
sorry, i forgot to multiply by 2.
so it's 1/8*2, since both sides count.
and equals to 1/4, which was your original answer :)
Join the discussion

by varnit » Sun Jul 20, 2008 10:42 pm
How I see it :

Either all three head or all three tail

(1/2)^3 + (1/2)^3
= 2/8 = 1/4
Join the discussion

by zagcollins » Mon Jul 21, 2008 4:24 am
not bad, varnit...thats a shortcut...thats what we need in math..faster ways to blow the problems out of the water!
Join the discussion

by Scott@TargetTestPrep » Wed Dec 13, 2017 1:11 pm
zagcollins wrote:When tossed, a certain coin has equal probability of landing on either side. If the coin is tossed 3 times, what is the probability that it will land on the same side each time?

A. 1/8
B. 1/4
C. 1/3
D. 3/8
E. 1/2
We need to determine the probability that a coin lands on the same side for all 3 tosses of the coin. This means that the 3-toss outcome will be either H-H-H or T-T-T.

P(T-T-T) = 1/2 x 1/2 x 1/2 = 1/8

P(H-H-H) = 1/2 x 1/2 x 1/2 = 1/8

Thus, the probability that the coin lands on the same side each time is 1/8 + 1/8 = 2/8 = 1/4.

Answer: B

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

by Brent@GMATPrepNow » Wed Dec 13, 2017 1:31 pm
zagcollins wrote:When tossed, a certain coin has equal probability of landing on either side. If the coin is tossed 3 times, what is the probability that it will land on the same side each time?

A. 1/8
B. 1/4
C. 1/3
D. 3/8
E. 1/2
P(3 matching tosses) = P(ANY result on 1st toss AND 2nd toss matches 1st toss AND 3rd toss matches 2nd toss)
= P(ANY result on 1st toss) x P(2nd toss matches 1st toss ) x P(3rd toss matches 2nd toss)
= 1 x 1/2 x 1/2
= 1/4
= B

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion